OpenStudy (anonymous):

Solve the triangle A=50 degrees, b=13, c=6 a=10.2,C=26.3,B=103.7 a=14,C=26.3,B=103.7 No triangles possible a=14,C=30.3,B=99.7

OpenStudy (anonymous):

law of sines

OpenStudy (anonymous):

I think you'll need law of cosines with that information.

OpenStudy (anonymous):

the question asks for law of sines

OpenStudy (anonymous):

Law of sines looks like this \[\large \frac{a}{sin(A)}=\frac{b}{sin(B)}=\frac{c}{sin(C)}\] Right now, all you have is \[\large \frac{a}{sin(50º)}=\frac{13}{sin(B)}=\frac{6}{sin(C)}\]

OpenStudy (anonymous):

It's possible with a little algebraic juggling, but I still think it's easier to use law of cosines first to get the third side, then law of sines will clean the rest up nicely.

OpenStudy (anonymous):

The law of cosines is: \(c^2 = a^2 + b^2 + 2ab \cdot cos C\) Other equations: \(a^2 = b^2 + c^2 - 2bc \cdot cos A\) <--- This one is the one you probably want. \(b^2 = a^2 + c^2 - 2ac \cdot cos B\) http://en.wikipedia.org/wiki/Law_of_cosines

OpenStudy (anonymous):

If you try it with law of sines alone, you'll need to set up a system of three equations for the three unknowns, and that will get a little annoying.

OpenStudy (anonymous):

i have to choose from the options above

OpenStudy (anonymous):

I still recommend finding a first, then you can use law of sines to get B and C.

OpenStudy (anonymous):

okay thankyou

OpenStudy (anonymous):

You're welcome.